a sum of rs 7700 is divided among three brothers zain ,zaid and zohaib in such a way that simple intrest on each part at 5% per annum after 1,2and 3 years respectively remains equal the share of zain is more than that of zohaib by
a) Rs 2,800 b)Rs 2500
c)Rs 3000 d)none of these

Answer by ankor@dixie-net.com(22740)

a sum of rs 7700 is divided among three brothers zain ,zaid and zohaib in such a way that simple intrest on each part at 5% per annum after 1,2and 3 years respectively remains equal
the share of zain is more than that of zohaib by ??
:
let's call these guy's shares a,b,c
a + b + c = 7700
:
"simple intrest on each part at 5% per annum after 1,2and 3 years respectively remains equal"
.05a = 2(.05b) = 3(.05c)
.05a = .10b = .15c
Get rid of the decimal, multiply by 100
5a = 10b = 15c
Find each in terms of a
10b = 5a
b = a/2
and
15c = 5a
c = a/3
therefore
a + a/2 + a/3 = 7700
Get rid of the fractions, multiply by 6
6a + 3a + 2a = 46200
11a = 46200
a = 46200/11
a = 4200
"the share of zain is more than that of zohaib by"
a - c = ??
Find c
c = 4200/3
c = 1400
Therefore 4200 - 1400 = 2800 which is a